# The spectrum of Dirichlet-to-Neumann maps for radial conductivities

### v1

@@ 2.2. The Born approximation for Schrödinger operator DtN maps
Let $d\geq 2$ and let $\gamma\in W^{2,p}_{\mathrm{rad}}(\mathbb{B}^{d})$ with $d/2<p\leq\infty$ be a conductivity. Let $V$ be given by (2.5) and $\Lambda^{\mathrm{p}}_{V}$ by (2.6). Then $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$ and there exists a unique $V^{\mathrm{B}}\in L^{1}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$ such that
(2.8) $\sigma_{k}[V^{\mathrm{B}}]=\lambda_{k}[V]-k,\qquad\text{for all }\,k\in\mathbb{N}_{0}.$
@@
    (2.9) $V^{\mathrm{B}}(x)=\frac{a_{0}}{|x|^{2}}+\sum_{j=1}^{J}\frac{a_{j}}{|x|^{2\kappa_{j}+2}}+V^{\mathrm{B}}_{\mathrm{reg}}(x),$
@@
    $V^{\mathrm{B}}|_{\mathbb{B}^{d}\setminus\{0\}}\in L^{p}_{\mathrm{loc}}(\mathbb{B}^{d}\setminus\{0\})$. In fact $V^{\mathrm{B}}-V$ is continuous on $\overline{\mathbb{B}^{d}}\setminus\{0\}$ and vanishes in $\partial\mathbb{B}^{d}$. Also, $V\in\mathcal{C}^{m}(U_{s})$ iff $V^{\mathrm{B}}\in\mathcal{C}^{m}(U_{s})$, for all $m\in\mathbb{N}_{0}$ and for all $0<s<1$.
Proof of Proposition˜2.2.
If $V$ is given by (2.5), then $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$, since $\gamma(x)>c>0$ in $\mathbb{B}^{d}$. From Lemma˜3.9 and Proposition˜3.10 i) it follows there exists a solution $V^{\mathrm{B}}$ of the moment problem (2.8). This solution is unique as follows from [BCMM22, Equation 1.20], which is reproduced here in (2.14) below. Then i) and ii) are immediate from Lemma˜3.9 and Proposition˜3.10. Finally, assertion iii) is proved in [DMMN24, Theorem 5]. ∎
@@ 2.3. The Born approximation for $\Lambda_{\gamma}$
The proof of ˜1 follows from Proposition˜2.2, Proposition˜2.1 and the following technical properties of radial functions.
@@
i) Since $\Delta u=f$, $f\in L^{p}(\mathbb{B}^{d})$ and $u$ vanishes on $\mathbb{S}^{d-1}$, $u\in W^{2,p}(\mathbb{B}^{d})$ by the standard Calderon-Zygmund estimates for the Laplacian.
ii) Let $F$ denote the radial profile of $u$. For all $\alpha>1$, we have the estimate
$|F^{\prime}(r)|\leq\frac{1}{r}\int_{0}^{r}|f_{0}(s)|s\,ds\leq\frac{1}{r(1-\log r)^{\alpha}}\int_{0}^{r}|f_{0}(s)|s(1-\log s)^{\alpha}\,ds\leq\frac{C_{\alpha}}{r(1-\log r)^{\alpha}}.$
This implies that $F(0)$ is well defined and finite, since, under our hypothesis on $\alpha$,
$|F(0)|\leq\int_{0}^{1}|F^{\prime}(r)|\,dr\leq\int_{0}^{1}\frac{1}{r(1-\log r)^{\alpha}}\,dr=\int_{0}^{\infty}\frac{1}{(1+t)^{\alpha}}dt<\infty.$
The integrability of $F^{\prime}$ in $[0,1]$ implies that $F$ is continuous on $[0,1]$, and hence, $u\in\mathcal{C}(\overline{\mathbb{B}^{d}})$. Since $\alpha>1>1/d$, one finds that the gradient satisfies
$\left\lVert\nabla u\right\rVert_{L^{d}(\mathbb{B}^{d})}^{d}\leq|\mathbb{S}^{d-1}|\int_{0}^{1}|F^{\prime}(r)|^{d}r^{d-1}\,dr\leq|\mathbb{S}^{d-1}|\int_{0}^{1}\frac{1}{r(1-\log r)^{d\alpha}}\,dr<\infty.$
@@
$\left\lVert D^{2}u\right\rVert_{L^{p}(\mathbb{B}^{d})}\leq(d+1)\left(\int_{\mathbb{B}^{d}}|F^{\prime}(|x|)|^{p}|x|^{d-1-p}\,dx\right)^{1/p}+\left\lVert f\right\rVert_{L^{p}(\mathbb{B}^{d})}\\ \leq C_{d,\alpha}\left(\int_{0}^{1}\frac{r^{d-1-2p}}{(1-\log r)^{p\alpha}}\,dr\right)^{1/p}+\left\lVert f\right\rVert_{L^{p}(\mathbb{B}^{d})}.$
The first integral converges whenever $1\leq p\leq d/2$, $\alpha>1$, and, by hypothesis, $\left\lVert f\right\rVert_{L^{1}(\mathbb{B}^{d})}<\infty$; therefore $\left\lVert D^{2}u\right\rVert_{L^{1}(\mathbb{B}^{d})}<\infty$ as we wanted to prove. ∎
@@
Let $V$ be given by (2.4) and restrict the operator identity (2.6) to functions in $\mathfrak{H}_{k}$ to obtain
(2.16) $\lambda^{\mathrm{p}}_{k}[V]-k=\frac{1}{a(\gamma)}\left(\lambda_{k}[\gamma]+b(\gamma)-a(\gamma)k\right),\qquad\forall\,k\in\mathbb{N}_{0},$
where the notation in (2.10) has been used. On the other hand, since $\gamma\in W^{2,p}_{\mathrm{rad}}(\mathbb{B}^{d})$, we have $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$. Therefore, Proposition˜2.2 applies and there exists $V^{\mathrm{B}}\in L^{1}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$ such that
$\lambda^{\mathrm{p}}_{k}[V]-k=\sigma_{k}[V^{\mathrm{B}}],\qquad\forall\,k\in\mathbb{N}_{0}.$
@@
Define
(2.18) $\gamma^{\mathrm{B}}(x):=a(\gamma)-2a(\gamma)T_{d}(V^{\mathrm{B}}).$
Then $\gamma^{\mathrm{B}}\in\mathcal{C}^{1}(\overline{\mathbb{B}^{d}}\setminus\{0\})$. By Proposition˜2.2, $V^{\mathrm{B}}=V^{\mathrm{B}}_{\mathrm{sing}}+V^{\mathrm{B}}_{\mathrm{reg}}$ where $\left\lVert V^{\mathrm{B}}_{\mathrm{reg}}\right\rVert_{d,\alpha}<\infty$ for all $\alpha>1$, and either $V^{\mathrm{B}}_{\mathrm{sing}}=0$, if $d=2$, or $V^{\mathrm{B}}_{\mathrm{sing}}\in L^{q}_{\mathrm{rad}}(\mathbb{B}^{d})$ for every $1\leq q<d/s_{V}$, with $s_{V}=2\max_{j=0,...,J}\{1,\kappa_{j}+1\}<d$ when $d\geq 3$. Therefore, by Lemma˜2.4, one gets that $\gamma^{\mathrm{B}}\in W^{2,1}_{\mathrm{rad}}(\mathbb{B}^{d})$. Also, $V^{\mathrm{B}}$ solves the Poisson problem
@@
(2.20) $\gamma_{1}|_{\mathbb{S}_{d-1}}=\gamma_{2}|_{\mathbb{S}_{d-1}},\qquad\qquad\partial_{\nu}\gamma_{1}|_{\mathbb{S}_{d-1}}=\partial_{\nu}\gamma_{2}|_{\mathbb{S}_{d-1}}.$
@@
Proof of ˜3.
This result is a straightforward consequence of ˜1, ˜2 and the following fact.
@@
One implication is straightforward; in order to check the other, assume that $|\sigma_{k}[F]|\leq Cs^{2k}$ for $k\geq k_{0}$. If $F=f(|\cdot|)$ one has $\sigma_{k}[F]=\mathcal{L}(F)(k+d/2)$ where
$\mathcal{L}(F)(z):=\int_{0}^{\infty}f(e^{-t})e^{-z2t}dt=\frac{1}{2}\int_{0}^{\infty}f(e^{-t/2})e^{-zt}dt.$
By hypothesis, $e^{(2k+d)|\log s|}|\mathcal{L}(F)(k+d/2)|\leq Ce^{-d|\log s|}$ for every $k\geq k_{0}$. Note that there exists $C_{s}>0$ such that
@@
$G(z):=e^{z2|\log s|}\int_{0}^{2|\log s|}f(e^{-t/2})e^{-zt}dt,$
we have, in view of (2.21), that $G$ is an entire function that satisfies
$\sup_{k\in\mathbb{N}}|G(k+d/2)|\leq Ce^{-d|\log s|}+C_{s}\left\lVert F\right\rVert_{L^{1}(\mathbb{B}^{d})}$
and
$|G(z)|\leq\left(\int_{0}^{2|\log s|}|f(e^{-t})|dt\right)e^{2|\log s|(\operatorname{\mathrm{Re}}z)_{+}}.$
@@
Using [Sim99, Lemma A.2.1] we infer that $f(t)=0$ for a.e. $t\in(0,s)$, and the conclusion follows. ∎
To see that two conductivities satisfying (1.9) must necessarily coincide on $U_{s}$, start noticing that (1.4) and (1.9) imply that $\gamma_{1}|_{\mathbb{S}^{d-1}}=\gamma_{2}|_{\mathbb{S}^{d-1}}$ and $\partial_{\nu}\gamma_{1}|_{\mathbb{S}^{d-1}}=\partial_{\nu}\gamma_{2}|_{\mathbb{S}^{d-1}}$. Lemma˜2.5 ensures then that $(\Delta\gamma^{\mathrm{B}}_{1})|_{U_{s}}=(\Delta\gamma^{\mathrm{B}}_{2})|_{U_{s}}$; therefore, by (1.5), $w:=\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}$ solves $\Delta w=0$ on $U_{s}$ with $w|_{\mathbb{S}^{d-1}}=\partial_{\nu}w|_{\mathbb{S}^{d-1}}=0$. This implies $w|_{U_{s}}=0$; the result follows now from ˜2. ∎
@@
By (2.18) we have that $\gamma^{\mathrm{B}}=a(\gamma)-2a(\gamma)T_{d}(V^{\mathrm{B}})$. Using this, the case ii) ($d=2$) and the regularity of $\gamma^{\mathrm{B}}_{\mathrm{reg}}$ in i) ($d\geq 3$) are immediate by Proposition˜2.2 and Lemma˜2.4 ii).
To finish the proof of i) notice that $T_{d}(|\cdot|^{-2\kappa-2})=\frac{r^{-2\kappa}-1}{2\kappa(d-2-2\kappa)}$ for all $0<\kappa<\nu_{d}$, and $T_{d}(|\cdot|^{-2})=-(d-2)^{-1}\log|x|$, $d\geq 3$. This yields using Proposition˜2.2 i) the representation in $d\geq 3$ with $c_{j}=-a_{j}\frac{1}{2\kappa(d-2-2\kappa_{j})}$ for $j=1,\dots,J$. Since $a_{j}\leq 0$ for all $j\geq 1$, then $c_{j}\geq 0$ for all $j\geq 1$. Finally iii) follows immediately from Proposition˜2.2 iii) using that $\gamma^{\mathrm{B}}$ and $\sqrt{\gamma}$ are, respectively, solutions of the elliptic equations $\Delta\gamma^{\mathrm{B}}=2a(\gamma)V^{\mathrm{B}}$ and $\Delta\sqrt{\gamma}=V\sqrt{\gamma}$. ∎

### v2

@@ 2.1. The Born approximation for DtN maps of Schrödinger operators
Let $d\geq 2$ and let $\gamma\in W^{2,p}_{\mathrm{rad}}(\mathbb{B}^{d})$ with $d/2<p\leq\infty$ be a conductivity. Let $V$ be given by (2.2) and $\Lambda^{\mathrm{S}}_{V}$ by (2.3). Then $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$ and there exists a unique $V^{\mathrm{B}}\in L^{1}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$ such that
(2.4) $\sigma_{k}[V^{\mathrm{B}}]=\lambda^{\mathrm{S}}_{k}[V]-k,\qquad\text{for all }\,k\in\mathbb{N}_{0}.$
@@
    (2.5) $V^{\mathrm{B}}(x)=V^{\mathrm{B}}_{\mathrm{reg}}(x)+\frac{a_{0}}{|x|^{2}}+\sum_{j=1}^{J}\frac{a_{j}}{|x|^{2\kappa_{j}+2}},$
@@
    $V^{\mathrm{B}}|_{\mathbb{B}^{d}\setminus\{0\}}\in L^{p}_{\mathrm{loc}}(\mathbb{B}^{d}\setminus\{0\})$. In fact, $V^{\mathrm{B}}-V$ is continuous on $\overline{\mathbb{B}^{d}}\setminus\{0\}$ and vanishes on $\partial\mathbb{B}^{d}$. Also, $V\in\mathcal{C}^{m}(U_{s})$ iff $V^{\mathrm{B}}\in\mathcal{C}^{m}(U_{s})$, for all $m\in\mathbb{N}_{0}$ and for all $0<s<1$.
  4. iv)
    $V^{\mathrm{B}}\in L^{1}(\mathbb{B}^{d})$.
Following [DMMN24], we say that $V^{\mathrm{B}}$ is the Born approximation of the potential $V$. Proposition˜2.1 is proved in Section˜3.4. It relies on a fine analysis of the singularities of the Born approximation and constitutes an important ingredient of the proof of ˜1 presented below.
@@ 2.2. The Born approximation for $\Lambda_{\gamma}$
The proof of ˜1 follows from Proposition˜2.1 and the following technical properties of radial functions.
@@
i) Since $\Delta u=f$, $f\in L^{p}(\mathbb{B}^{d})$ and $u$ vanishes on $\mathbb{S}^{d-1}$, $u\in W^{2,p}(\mathbb{B}^{d})$ by the standard Calderón-Zygmund estimates for the Laplacian.
ii) Denote by $F$ and $f_{0}$, respectively, the radial profiles of $u=T_{d}(f)$ and $f$. By (2.11) we have
$|F^{\prime}(r)|\leq\frac{1}{r^{d-1}}\int_{0}^{r}|f_{0}(s)|s^{d-1}\,ds\leq\frac{1}{r}\int_{0}^{r}|f_{0}(s)|s\,ds.$
Then, for all $\alpha>1$, using that $(1-\log r)\leq(1-\log s)$ for $s<r$, we have the estimate
$|F^{\prime}(r)|\leq\frac{1}{r(1-\log r)^{\alpha}}\int_{0}^{r}|f_{0}(s)|s(1-\log s)^{\alpha}\,ds\leq\frac{c_{d}\left\lVert f\right\rVert_{d,\alpha}}{r(1-\log r)^{\alpha}},$
where $c_{d}:=|\mathbb{S}^{d-1}|^{-1}$. This implies that $F(0)$ is well defined and finite, since, under our hypothesis on $\alpha$,
$|F(0)|\leq\int_{0}^{1}|F^{\prime}(r)|\,dr\leq\int_{0}^{1}\frac{c_{d}\left\lVert f\right\rVert_{d,\alpha}}{r(1-\log r)^{\alpha}}\,dr=\int_{0}^{\infty}\frac{c_{d}\left\lVert f\right\rVert_{d,\alpha}}{(1+t)^{\alpha}}dt<\infty.$
The integrability of $F^{\prime}$ in $[0,1]$ implies that $F$ is continuous on $[0,1]$, and hence, $u\in\mathcal{C}(\overline{\mathbb{B}^{d}})$. Since $\alpha>1$ then $\alpha d>1$ and one finds that the gradient satisfies
$\left\lVert\nabla u\right\rVert_{L^{d}(\mathbb{B}^{d})}^{d}\leq|\mathbb{S}^{d-1}|\int_{0}^{1}|F^{\prime}(r)|^{d}r^{d-1}\,dr\leq\int_{0}^{1}\frac{c_{d}^{d-1}\left\lVert f\right\rVert_{d,\alpha}^{d}}{r(1-\log r)^{d\alpha}}\,dr<\infty.$
@@
$\left\lVert D^{2}u\right\rVert_{L^{1}(\mathbb{B}^{d})}$ $\leq(d+1)\int_{\mathbb{B}^{d}}|F^{\prime}(|x|)||x|^{-1}\,dx+\left\lVert f\right\rVert_{L^{1}(\mathbb{B}^{d})}$
$\leq(d+1)\left\lVert f\right\rVert_{d,\alpha}\int_{0}^{1}\frac{r^{d-3}}{(1-\log r)^{\alpha}}\,dr+\left\lVert f\right\rVert_{L^{1}(\mathbb{B}^{d})}.$
The first integral converges since $\alpha>1$, and, by hypothesis, $\left\lVert f\right\rVert_{L^{1}(\mathbb{B}^{d})}<\infty$; therefore $\left\lVert D^{2}u\right\rVert_{L^{1}(\mathbb{B}^{d})}<\infty$ as we wanted to prove. ∎
@@
Let $V$ be given by (2.2) and restrict the operator identity (2.3) to functions in $\mathfrak{H}_{k,d}$ to obtain
(2.12) $\lambda^{\mathrm{S}}_{k}[V]-k=\frac{1}{a(\gamma)}\left(\lambda_{k}[\gamma]+b(\gamma)-a(\gamma)k\right),\qquad\forall\,k\in\mathbb{N}_{0},$
where the notation in (2.6) has been used. On the other hand, since $\gamma\in W^{2,p}_{\mathrm{rad}}(\mathbb{B}^{d})$, we have $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$. Therefore, Proposition˜2.1 applies and there exists $V^{\mathrm{B}}\in L^{1}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$ such that
$\lambda^{\mathrm{S}}_{k}[V]-k=\sigma_{k}[V^{\mathrm{B}}],\qquad\forall\,k\in\mathbb{N}_{0}.$
@@
We now choose $\gamma^{\mathrm{B}}$ so that $\Delta\gamma^{\mathrm{B}}=2a(\gamma)V^{\mathrm{B}}$. In order to do so, define
(2.14) $\gamma^{\mathrm{B}}(x):=a(\gamma)-2a(\gamma)T_{d}(V^{\mathrm{B}}),$
where $T_{d}$ was defined in (2.11). Then $\gamma^{\mathrm{B}}\in\mathcal{C}^{1}(\overline{\mathbb{B}^{d}}\setminus\{0\})$.
By Proposition˜2.1, $V^{\mathrm{B}}=V^{\mathrm{B}}_{\mathrm{sing}}+V^{\mathrm{B}}_{\mathrm{reg}}$ where $\left\lVert V^{\mathrm{B}}_{\mathrm{reg}}\right\rVert_{d,\alpha}<\infty$ for all $\alpha>1$, and either $V^{\mathrm{B}}_{\mathrm{sing}}=0$, if $d=2$, or $V^{\mathrm{B}}_{\mathrm{sing}}\in L^{q}_{\mathrm{rad}}(\mathbb{B}^{d})$ for every $1\leq q<d/s_{V}$, with $s_{V}=2\max_{j=0,...,J}\{1,\kappa_{j}+1\}<d$ when $d\geq 3$. Therefore, by Lemma˜2.3, one gets that $\gamma^{\mathrm{B}}\in W^{2,1}_{\mathrm{rad}}(\mathbb{B}^{d})$. Also, $\gamma^{\mathrm{B}}$ solves the Poisson problem
@@ 2.3. Proof of the local uniqueness results
(2.16) $\gamma_{1}|_{\mathbb{S}^{d-1}}=\gamma_{2}|_{\mathbb{S}^{d-1}},\qquad\qquad\partial_{\nu}\gamma_{1}|_{\mathbb{S}^{d-1}}=\partial_{\nu}\gamma_{2}|_{\mathbb{S}^{d-1}}.$
@@
Before proving ˜3, we need to prove a local uniqueness result for the Hausdorff Moment Problem.
@@
If $F|_{U_{s}}=0$ it is straightforward to verify that $|\sigma_{k}[F]|\leq Cs^{2k}$ for all $k\geq 0$. It remains to check the other implication. If $F=f(|\cdot|)$ one has $\sigma_{k}[F]=\mathcal{L}(F)(k+d/2)$ where
$\mathcal{L}(F)(z):=\int_{0}^{\infty}f(e^{-t})e^{-z2t}dt=\frac{1}{2}\int_{0}^{\infty}f(e^{-t/2})e^{-zt}dt,$
and there exists $C_{s}>0$ such that
@@
$G(z):=e^{z2|\log s|}\int_{0}^{2|\log s|}f(e^{-t/2})e^{-zt}dt.$
By hypothesis, $|\sigma_{k}[F]|\leq Cs^{2k}$ for $k\geq k_{0}$, which implies that
$e^{(2k+d)|\log s|}|\mathcal{L}(F)(k+d/2)|\leq Cs^{-d},\qquad\forall k\geq k_{0}.$
Hence, in view of (2.17), $G$ satisfies
$|G(k+d/2)|\leq Cs^{-d}+C_{s}\left\lVert F\right\rVert_{L^{1}(\mathbb{B}^{d})}\qquad\forall k\geq k_{0}.$
Thus, $G$ is an entire function that satisfies $\sup_{k\in\mathbb{N}}|G(k+d/2)|\leq\tilde{C}$ for some $\tilde{C}>0$ and
$|G(z)|\leq\left(\int_{0}^{2|\log s|}|f(e^{-t/2})|dt\right)e^{2|\log s|(\operatorname{\mathrm{Re}}z)_{+}}.$
@@
Using the local injectivity of the Laplace transform, see [Sim99, Lemma A.2.1], we infer that $f(e^{-t/2})=0$ for a.e. $t\in(0,2|\log s|)$, hence $f(r)=0$ for a.e. $r\in(s,1)$, and the conclusion follows. ∎
Proof of ˜3.
This result is a straightforward consequence of ˜1, ˜2 and Lemma˜2.4. To see that two conductivities satisfying (1.9) must necessarily coincide on $U_{s}$, start by noticing that (1.4) and (1.9) imply that $\gamma_{1}|_{\mathbb{S}^{d-1}}=\gamma_{2}|_{\mathbb{S}^{d-1}}$ and $\partial_{\nu}\gamma_{1}|_{\mathbb{S}^{d-1}}=\partial_{\nu}\gamma_{2}|_{\mathbb{S}^{d-1}}$. Lemma˜2.4 ensures then that $(\Delta\gamma^{\mathrm{B}}_{1})|_{U_{s}}=(\Delta\gamma^{\mathrm{B}}_{2})|_{U_{s}}$; therefore, by (1.5), $w:=\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}$ solves $\Delta w=0$ on $U_{s}$ with $w|_{\mathbb{S}^{d-1}}=\partial_{\nu}w|_{\mathbb{S}^{d-1}}=0$. This implies $w|_{U_{s}}=0$; the result follows now from ˜2. ∎
@@
By (2.14) we have that $\gamma^{\mathrm{B}}=a(\gamma)-2a(\gamma)T_{d}(V^{\mathrm{B}})$. Using this, the case ii) ($d=2$) and the regularity of $\gamma^{\mathrm{B}}_{\mathrm{reg}}$ in i) ($d\geq 3$) are immediate by Proposition˜2.1 and Lemma˜2.3 ii).
To finish the proof of i) notice that $T_{d}(|\cdot|^{-2\kappa-2})=\frac{|x|^{-2\kappa}-1}{2\kappa(d-2-2\kappa)}$ for all $0<\kappa<\nu_{d}$, and $T_{d}(|\cdot|^{-2})=-(d-2)^{-1}\log|x|$, $d\geq 3$. This yields using Proposition˜2.1 i) the representation in $d\geq 3$ with $c_{j}=-a_{j}\frac{1}{2\kappa_{j}(d-2-2\kappa_{j})}$ for $j=1,\dots,J$. Since $a_{j}\leq 0$ for all $j\geq 1$, then $c_{j}\geq 0$ for all $j\geq 1$. Finally iii) follows immediately from Proposition˜2.1 iii) using that $\gamma^{\mathrm{B}}$ and $\sqrt{\gamma}$ are, respectively, solutions of the elliptic equations $\Delta\gamma^{\mathrm{B}}=2a(\gamma)V^{\mathrm{B}}$ and $\Delta\sqrt{\gamma}=V\sqrt{\gamma}$. ∎